14(x-4)-(x+2)=(x+2)(x-4)
Answer by stanbon(75887)

14(x-4)-(x+2)=(x+2)(x-4)
14x-56-x-2 = x^2-2x-8
13x-58 = x^2-2x-8
x^2-15x+50=0
(x-10)(x-5)=0
x = 10 or x = 5
=====================
Cheers,
Stan H.